OpenStudy (anonymous):

What is the slope of the line passing through the points (0, 4) and (−8, −1)?

OpenStudy (anonymous):

@triciaal

OpenStudy (triciaal):

follow the same steps m = (y2 - y1)/(x2 - x1) points( x1, y1) and (x2, y2)

OpenStudy (triciaal):

what do you get ?

OpenStudy (triciaal):

m =(-1-4)/(-8 - 0) = -5/-8= 5/8

OpenStudy (triciaal):

negative divided by negative gives positive